The mean monthly expenditure on gasoline per household in Middletown is determined by selecting a random sample of 100 households

24.

The mean monthly expenditure on gasoline per household in Middletown is determined by selecting a random sample of 100 households. The sample mean is $115, with a sample standard deviation of $38. What is the upper bound of a 80% confidence interval for the mean monthly expenditure on gasoline per household in Middletown?

$118.42

$116.38

$121.66

- $149.20

$119.87


25.

You estimated the average rental cost of a 2 bedroom apartment in Denton. A random sample of 16 apartments was taken. The sample mean is $ 800 with a known standard deviation of $ 160. The Error for this calculation, using a 95 % Cl is $78.40. You decide you cannot have this large an error and want to reduce the error to $ 60. What size sample should you take?

28

25

37

33 .

40
Answer #1

24:

Here we have

n=100,\bar{x}=115,s=38

Degree of freedom: df=n-1=37

The critica value for upper tail 80% confidence interval using excel "=TINV(0.20,99)" is 1.290.

The upper bound is

\bar{x}+t_{c}\cdot \frac{s}{\sqrt{n}}=115+1.29\cdot \frac{38}{\sqrt{100}}=119.902

Last option is correct.